What’s the measure of all the angles of the triangles if the ratio of the measure of the angles is 3:6:1?
Answer by om(1)

let the angle be 3x,6x,x
3x+6x+x=180 (sum of angles of triangle)
10x=180
x=180/10
x=18

.
. . x=18
3x=18*3=54
6x=18*6=108